OpenStudy (anonymous):

A 50-N ⋅ m torque acts on a wheel with a moment of inertia 150 kg ⋅ m2. If the wheel starts from rest, how long will it take the wheel to make one revolution? (a) 0.33 s (b) 2.4 s (c) 10 s (d) 0.66 s (d) 6.1 s

OpenStudy (vincent-lyon.fr):

This is basic application of kinematics concerning rotation and angular acceleration. What do you think?

OpenStudy (anonymous):

im not sure of the formula to use for angular acceleration?

OpenStudy (anonymous):

torque=MOMENT OF INERTIA*angular acceleration

OpenStudy (vincent-lyon.fr):

The formulae are the same as those for linear acceleration: Substitute: Torque τ for Force F Moment of inertia I for Mass m Angular acceleration α for Linear acceleration a Then substitute: Angular velocity ω for Linear velocity v Angular displacement θ for Linear displacement s (or x)

OpenStudy (anonymous):

Remember formula: \tau = I*\alpha. Solve for \alpha. Then 2*\pi = 1/2*\alpha*t^2 and solve for t.